OpenStudy (anonymous):

a jar contains 4 green marbles, 5 red marbles, and 7 blue marbles. If two balls are selected randomly what is the probability of getting : 1. 1 green and 1 blue 2. 2 red 3. 1 blue and 1 red 4. 2 blue

OpenStudy (anonymous):

hey help please?

OpenStudy (yrelhan4):

well i just got confused.. give me a minute.

OpenStudy (anonymous):

okay the title of lesson is combinations i got no idea how to do it lol

OpenStudy (anonymous):

@ganeshie8

OpenStudy (zarkon):

are you sampling with or without replacement?

OpenStudy (anonymous):

without i think

OpenStudy (zarkon):

then this is a hypergeometric distribution

OpenStudy (anonymous):

a what????? lol explain

OpenStudy (zarkon):

suppose you have a box with a red and b green balls. you grab k balls the prob you get r red and g green balls (r+g=k) is given by \[\frac{ _aC_r _bC_g}{_{(a+b)}C_k}\]

OpenStudy (anonymous):

yeah i dont get it

OpenStudy (zarkon):

\[\frac{_aC_r\cdot_b C_g}{_{a+b}C_k}\]

OpenStudy (zarkon):

do you know what \[_nC_r\] is ?

OpenStudy (anonymous):

not at all

OpenStudy (zarkon):

\[{n\choose r}\]?

OpenStudy (anonymous):

nope i dont know any of this stuff

OpenStudy (zarkon):

permutations/combinations?

OpenStudy (yrelhan4):

i think it would be easier if we dont use combinations, zarkon.. it would be easier if if we do it like this.. for 1 green ball. 4/16.. 1 blue ball after 1 green ball hass been selected.. 7/15. and then multiply them.. whats your view?

OpenStudy (anonymous):

i knwo thats what this unit is called combinations

OpenStudy (yrelhan4):

eh right. so it would be like 4/16*7/15 + 7/16*4/15.. right?

OpenStudy (anonymous):

id think s if that was the case

OpenStudy (anonymous):

wait but youd get the same thing right 7/60 + 7/60

OpenStudy (yrelhan4):

yup.

OpenStudy (yrelhan4):

then for the second part.. 5/16*4/15.. you get why?

OpenStudy (anonymous):

which would equal 14/60 right because you dont change the denominator right or would it be 14/120

OpenStudy (yrelhan4):

yeah.

OpenStudy (anonymous):

so it is 14/120

OpenStudy (yrelhan4):

am i right with the second part @RolyPoly ?

OpenStudy (yrelhan4):

it should be 14/120..

OpenStudy (zarkon):

no

OpenStudy (yrelhan4):

14/60.. lol

OpenStudy (zarkon):

the answer is 7/30

OpenStudy (anonymous):

well idk how to do this thn i am so lost im gonna fail lol

OpenStudy (zarkon):

or 14/60 if you like

OpenStudy (anonymous):

how

OpenStudy (zarkon):

the solution is written above

OpenStudy (anonymous):

okay so will you help by showing work on two three and four maybe if i see it on three ill be able to do it

OpenStudy (anonymous):

P(blue first, then green) = P(blue) x P(green after drawing a blue) = 7/16 x 4/(16-1) = a P(green first, then blue) = P(green) x P(blue after drawing a green) = 4/16 x 7/(16-1) = b P(1 blue and 1 green) =P(blue first, then green) + P(green first, then blue) = a + b

OpenStudy (anonymous):

now im deffinately confused i just ned a simple explanation

OpenStudy (anonymous):

2 red, similar, despite the order of picking a red doesn't matter this time, since you don't the red ones are identical

OpenStudy (anonymous):

what omg i need a tutor

OpenStudy (anonymous):

Which part are you confused at?

OpenStudy (anonymous):

2 BLUE

OpenStudy (anonymous):

2 red actually all of it really

OpenStudy (anonymous):

u gonna help k?

OpenStudy (anonymous):

no thats not the question @phi help please